0.1 mole nitrogen is heated from 27°C to 327°C at constant pressure. Determine the heat given to the gas, increase in internal energy and work done by the gas. For Nitrogen, cp= 7 and cv= 5 Cal mol-1 °C-¹.Please solve fast.. It`s an urgent and humble request.

Grade:11

2 Answers

Pranav Prakasan
6 years ago
The change in internal energy is = 600 J. The amount of heat energy provided is = 840 J.U = Q - WW = P V
Dipika Biswakarma
3 years ago
Heat given to gas, h = nCp∆T 
= 0.1 × 7×300°C 
= 210 cal 
Internal energy, U = nCv∆T 
      = 0.1 ×5×300°C 
      = 150 cal 
Work done by had , W = h+U 
= 210 + 150 
= 360 cal